OpenStudy (anonymous):

x+30=x under a square root

OpenStudy (nory):

What's under the square root? Do you mean \[\sqrt{x+30} = x\] or \[x + 30 = \sqrt{x}\] or what?

OpenStudy (anonymous):

the first one you just said

OpenStudy (espex):

How far have you gotten in your attempt to solve?

OpenStudy (anonymous):

no where

OpenStudy (espex):

Your first step is going to look at how you can get rid of the radical.

OpenStudy (anonymous):

I'm not sure how to do it. That's why I'm here.

OpenStudy (espex):

Start by looking at your left side, to get rid of your square root, you must square it, and what you do to one side, you must do to the other.

OpenStudy (strawberry17):

Like this: \[(\sqrt{x + 30})^{2} = (x)^{2}\] Then when you square both sides, you can take away the square root of the left side and you are left with: x + 30 = x^2

OpenStudy (strawberry17):

I mean take away the radical on the left side..

OpenStudy (nory):

Now you can solve the equation that @strawberry17 gave you by putting all the terms on one side, and because it's a quadratic equation, you can now easily solve it.

OpenStudy (nory):

Don't forget to plug in your solutions when you're done so that you make sure they work out in the original equation. Always important, especially with radical equations.

OpenStudy (nory):

No. Your equations were correct. I am telling WoozySlick how to finish up the equation.